OpenStudy (anonymous):

(2x-1)/(x+1) - (2x-1)/(x+2) ; I know how to solve it, but I keep getting the wrong answer

OpenStudy (anonymous):

why don't you show what you've tried, so we can correct you :) ?

OpenStudy (anonymous):

\[\LARGE {2x-1 \over x+1}-{2x-1 \over x+2}=\] \[\LARGE {(2x-1)(x+2) \over (x+1)(x+2)}-{(2x-1)(x+1) \over (x+2)(x+1)}=\] \[\LARGE {2x^2+4x-x-2 \over (x+1)(x+2)}-{2x^2+2x-x-1\over (x+2)(x+1)}=\] \[\LARGE {2x^2+3x -2 \over (x+1)(x+2)}-{2x^2+x-1\over (x+2)(x+1)}=\] \[\LARGE {2x^2+3x -2-(2x^2+x-1) \over (x+1)(x+2)}=\] \[\LARGE {2x^2+3x -2-2x^2-x+1 \over x^2+2x+x+2}=\] \[\LARGE {2x -1 \over x^2+3x+2}=\]

OpenStudy (anonymous):

I hope I didn't get the wrong answer too... lol :B

OpenStudy (anonymous):

That's what I got. haha but I actually made a mistake and solved the equation for x+2 instead of x-1

OpenStudy (anonymous):

x-2*..but thank you!